Determine the intensity of shear of an oil having viscosity = 1.2 poise and is used for lubrication in the clearance between a 10 cm diameter shaft and its journal bearing. The clearance is 1.0 mm and shaft rotates at 200 r.p.m.
1
Expert's answer
2020-11-10T20:01:54-0500

Given:

"\mu=1.2poise=0.12N\cdot s/m^2"

Diameter of shaft "D=10cm=0.1m"

Distance between shaft and journal bearing "dy=1.0mm=10^{-3}m"

Speed of shaft "N=200\ r.p.m."


Tangential speed of shaft is given by


"u=\dfrac{\pi DN}{60}=\dfrac{\pi\cdot0.1m\cdot200\ r.p.m.}{60}=\dfrac{\pi}{3}m/s"

Change of velocity between shaft and bearing


"du=u-0=\dfrac{\pi}{3}m/s"

Mathematically


"\tau=\mu\dfrac{du}{dy}=0.12N\cdot s/m^2\times\dfrac{\dfrac{\pi}{3}m/s}{10^{-3}m}="

"=40\pi N/m^2\approx125.6637 N/m^2"